# Euclid Quick Data Release (Q1)

### v1

@@ 5.1 Morphological properties of the bright quasar sample
The VIS PSF fraction ($f_{{\scriptscriptstyle\rm PSF}}$) provides a more direct handle on this unresolved light. Its distribution peaks around 0.8 with another weaker peak around 0.1 (Fig. 12), confirming that most sources are dominated by their cores, consistent with their Srsic fits converging to extreme values. Sources with high $f_{{\scriptscriptstyle\rm PSF}}$ values also correspond to the most point-like objects, characterised by $\mu_{\mathrm{max}}-\mathrm{mag}\approx-2.9$. However, low $f_{{\scriptscriptstyle\rm PSF}}$ values (e.g., $f_{{\scriptscriptstyle\rm PSF}}<0.1$) do not necessarily imply the absence of AGN activity, because uncertainties of the measurements could dominate at such low levels.
As shown in Fig. 12 and Table 4, the CAS parameters of the intermediate-redshift quasars show very different distributions compared to those of the low-redshift quasars. Highly left-skewed distributions of concentration (median $C=2.56$) and clumpiness (median $S=0.08$) are seen in intermediate-redshift quasars, in contrast to the nearly symmetric distributions of the two parameters for the low-redshift quasars (median $C=4.27$ and median $S=0.19$). The concentration and clumpiness distributions of the intermediate-redshift quasars appear unphysical, because quasar hosts are expected to be more concentrated and clumpier towards higher redshift.This discrepancy likely arises because the nuclear PSF components were not removed before the measurements. Indeed, sources with low nuclear fractions ($f_{\mathrm{PSF}}\lesssim 0.3$) show more realistic Srsic indices and CAS parameters, reinforcing that uncorrected nuclear light biases the structural measurements of the host galaxies.
Taken together, these results demonstrate that VIS morphology captures a strong redshift dependence in the apparent structure of bright quasars: (i) at $z<0.5$, obvious host galaxy structures are frequently resolved, and a single Srsic model can describe the light profiles of half of the sources; (ii) at $0.5<z<2$, the nuclear component dominates, driving Srsic indices of 90 % of the sources to their upper limits. In this regime ($0.5<z<2$), $f_{{\scriptscriptstyle\rm PSF}}$ offers a more reliable compactness measure and light profile characterisation than the single Srsic fits, allowing us to quantify the balance between nuclear and host emission across the bright quasar sample. Representative Euclid 1D spectra and VIS cutouts illustrating these trends are shown in Appendix C for low-redshift sources ($z<0.5$; Fig. 14) and for $0.5<z<2$ quasars spanning different $f_{{\scriptscriptstyle\rm PSF}}$ levels (Fig. 15).
@@ 5.2 Euclid colour space, multi-wavelength selection, and future improvements
The advent of Euclid’s high-quality, wide-area near-infrared photometry opens a novel colour space for quasar selection, providing new opportunities to discover dust-reddened or intrinsically red quasars that are missed by traditional optical or UV selection. The extended wavelength coverage of the Euclid NIR bands, combined with precise VIS photometry, enables more robust discrimination of quasars from stars and galaxies, particularly at redshifts and extinction levels where classical techniques are less effective. Our initial exploration demonstrates that red quasars occupy distinct loci in Euclid colour-colour diagrams, highlighting the potential of the survey to identify populations that are heavily underrepresented in previous optical surveys.
Mid-infrared photometry from WISE has already proven highly effective for selecting luminous AGN and quasars via their characteristic red MIR colours. Our analysis confirms that WISE-based colour cuts efficiently select AGN-dominated sources, as reflected in the properties of our pre-selected sample. However, WISE alone cannot distinguish between the most dust-obscured and less obscured quasars. The synergy between Euclid and WISE thus offers a powerful selection strategy: by combining WISE MIR colours with deep VIS and NIR photometry from Euclid, future quasar searches can achieve high completeness, efficiently probing obscured AGN populations.
Moreover, incorporating multi-band photometric data into the spectroscopic identification process will substantially improve the completeness and reliability of the quasar sample. Photometric pre-selection not only increases the likelihood of identifying genuine quasars, especially those with atypical spectral energy distributions, but also enhances redshift determination by providing prior constraints and mitigating degeneracies in the spectral fitting. The inclusion of Euclid and external photometry will be essential for robust identification and breaking redshift degeneracies from only one emission line.
A limitation of the present work is that the initial sample relies on pre-selection using Gaia and WISE catalogues, which may miss quasars outside the colour selection windows or at the limits of the Gaia and WISE sensitivity. In future Euclid data releases, the construction of a more complete and unbiased quasar sample will be enabled by combining traditional pre-selected candidates with new candidates identified from Euclid photometry and by the SPE pipeline. This composite approach, leveraging the full capability of the Euclid data, will allow for the recovery of quasars with a wider range of colours, luminosities, and host properties.
@@ 6 Conclusions
In this work, we have presented a large sample of approximately 3500 bright quasars covering the redshift range of $0<z\lesssim 4.8$ identified with red-grism spectroscopy of the first Euclid Quick Data Release. To ensure identification efficiency, we focus on a high-purity input quasar candidate sample of 9214 sources based on GDR3-QSOs (Quaia, CatNorth, and CatSouth) and the AllWISE R90 AGN table. Through a template matching process based on the Pearson correlation function and a visual inspection campaign, we label quasars and determine their spectroscopic redshifts. The success rate of the spectroscopic identification is 38 %.
From the identified sample, we estimate an empirical spectroscopic depth of $J_{\scriptscriptstyle\rm E}\lesssim 21.5$ and $H_{\scriptscriptstyle\rm E}\lesssim 21.3$ at the sensitivity of the Wide Field Survey, beyond which the number of securely identified quasars declines sharply. Our investigation of the novel Euclid colour space demonstrates its power in uncovering redder, dust-obscured quasars that may be missed by traditional optical and MIR selections. The synergy between Euclid and WISE photometry promises even greater completeness and diversity in future quasar surveys.
We construct the first Euclid composite spectrum of bright quasars, covering rest-frame NUV to NIR wavelengths and free from telluric absorption, providing a valuable benchmark for future spectral studies. The spectroscopic bright quasar catalogue of this work, and the composite quasar spectrum, will be available at https://cdsarc.cds.unistra.fr/. In addition, the spectral properties of the quasars will be derived by multi-component fitting (Euclid Collaboration: Calhau et al. in preparation).
We characterise VIS morphologies using single-Srsic and model-independent (CAS) parameters, supplemented by a deep-learning PSF fraction, $f_{{\scriptscriptstyle\rm PSF}}$. At low redshift ($z<0.5$), obvious host structures are frequently resolved, and a single Srsic model describes the light profiles of about half of the sources. At intermediate redshift ($0.5<z<2$), the nuclear component dominates, driving Srsic indices of roughly 90 % of the sources to the upper bound; in this regime $f_{{\scriptscriptstyle\rm PSF}}$ provides a more reliable compactness measure and is used to quantify the balance between nuclear and host emission across the sample.
Finally, we discuss current limitations, including the pre-selection bias and the challenges of host-nucleus decomposition, and outline pathways for improvement with upcoming Euclid data releases and advanced selection methods. The results presented here illustrate the transformative impact of Euclid on quasar science and set the stage for larger, more complete AGN catalogues in the era of wide-field space-based surveys.
@@ Acknowledgements.
The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMIMI, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org/consortium/community/). This work has made use of the Euclid Quick Release Q1 data from the Euclid mission of the European Space Agency (ESA), 2025, https://doi.org/10.57780/esa-2853f3b. This research makes use of ESA Datalabs (datalabs.esa.int), an initiative by ESAs Data Science and Archives Division in the Science and Operations Department, Directorate of Science. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, and NEOWISE, which is a project of the Jet Propulsion Laboratory/California Institute of Technology. WISE and NEOWISE are funded by the National Aeronautics and Space Administration. This publication makes use of data from the Hyper Suprime-Cam (HSC). This research made use of hips2fits [^6] https://alasky.cds.unistra.fr/hips-image-services/hips2fits, a service provided by CDS. YF is financed by the Dutch Research Council (NWO) grant number OCENW.XL21.XL21.025. KIC acknowledges funding from the Dutch Research Council (NWO) through the award of the Vici Grant VI.C.212.036. VA, JC and FR acknowledge the support from the INAF Large Grant ”AGN & Euclid: a close entanglement” Ob. Fu. 01.05.23.01.14.
Appendix A Description of the spectroscopically identified bright quasar catalogue

### v2

@@ 5.1 Morphological properties of the bright quasar sample
The VIS PSF fraction ($f_{{\scriptscriptstyle\rm PSF}}$) provides a more direct handle on this unresolved light. Its distribution peaks around 0.8 with another weaker peak around 0.1 (Fig. 12), confirming that most sources are dominated by their cores, consistent with their Sérsic fits converging to extreme values. Sources with high $f_{{\scriptscriptstyle\rm PSF}}$ values also correspond to the most point-like objects, characterised by $\mu_{\mathrm{max}}-\mathrm{mag}\approx-2.9$. However, low $f_{{\scriptscriptstyle\rm PSF}}$ values (e.g. $f_{{\scriptscriptstyle\rm PSF}}<0.1$) do not necessarily imply the absence of AGN activity, because uncertainties of the measurements could dominate at such low levels.
As shown in Fig. 12 and Table 4, the CAS parameters of the intermediate-redshift quasars show very different distributions compared to those of the low-redshift quasars. Highly left-skewed distributions of concentration (median $C=2.56$) and clumpiness (median $S=0.08$) are seen in intermediate-redshift quasars, in contrast to the nearly symmetric distributions of the two parameters for the low-redshift quasars (median $C=4.27$ and median $S=0.19$). The concentration and clumpiness distributions of the intermediate-redshift quasars appear unphysical, because quasar hosts are expected to be more concentrated and clumpier towards higher redshift.This discrepancy likely arises because the nuclear PSF components were not removed before the measurements. Indeed, sources with low nuclear fractions ($f_{\mathrm{PSF}}\lesssim 0.3$) show more realistic Sérsic indices and CAS parameters, reinforcing that uncorrected nuclear light biases the structural measurements of the host galaxies.
Taken together, these results demonstrate that VIS morphology captures a strong redshift dependence in the apparent structure of bright quasars: (i) at $z<0.5$, obvious host galaxy structures are frequently resolved, and a single Sérsic model can describe the light profiles of half of the sources; (ii) at $0.5<z<2$, the nuclear component dominates, driving Sérsic indices of 90 % of the sources to their upper limits. In this regime ($0.5<z<2$), $f_{{\scriptscriptstyle\rm PSF}}$ offers a more reliable compactness measure and light profile characterisation than the single Sérsic fits, allowing us to quantify the balance between nuclear and host emission across the bright quasar sample. Representative Euclid 1D spectra and VIS cutouts illustrating these trends are shown in Appendix D for low-redshift sources ($z<0.5$; Fig. 15) and for $0.5<z<2$ quasars spanning different $f_{{\scriptscriptstyle\rm PSF}}$ levels (Fig. 16). A more detailed analysis of AGN host galaxy morphology, including re-measurement of morphological parameters with methods tailored to AGN host galaxies, and dedicated tests to quantify measurement sensitivity as a function of host galaxy surface brightness, redshift, and nucleus-to-host contrast, will be presented in a follow-up study.
@@ 5.2 Euclid colour space, multi-wavelength selection, and future improvements
The advent of Euclid’s high-quality, wide-area near-infrared photometry opens a novel colour space for quasar selection, providing new opportunities to discover dust-reddened or intrinsically red quasars that are missed by traditional optical or ultraviolet (UV) selection. The extended wavelength coverage of the Euclid NIR bands, combined with precise VIS photometry, enables more robust discrimination of quasars from stars and galaxies, particularly at redshifts and extinction levels where classical techniques are less effective. Our initial exploration demonstrates that red quasars occupy distinct loci in Euclid colour-colour diagrams, highlighting the potential of the survey to identify populations that are heavily underrepresented in previous optical surveys.
Mid-infrared photometry from WISE has already proven highly effective for selecting luminous AGNs and quasars via their characteristic red MIR colours. Our analysis confirms that WISE-based colour cuts efficiently select AGN-dominated sources, as reflected in the properties of our pre-selected sample. However, WISE alone cannot distinguish between the most dust-obscured and less obscured quasars. The synergy between Euclid and WISE thus offers a powerful selection strategy: by combining WISE MIR colours with deep VIS and NIR photometry from Euclid, future quasar searches can achieve high completeness, efficiently probing obscured AGN populations.
Moreover, incorporating multi-band photometric data into the spectroscopic identification process will substantially improve the completeness and reliability of the quasar sample. Photometric pre-selection not only increases the likelihood of identifying genuine quasars, especially those with atypical SEDs, but also enhances redshift determination by providing prior constraints and mitigating degeneracies in the spectral fitting. The inclusion of Euclid and external photometry will be essential for robust identification and breaking redshift degeneracies from only one emission line.
A limitation of the present work is that the initial sample relies on pre-selection using Gaia and WISE catalogues, which may miss quasars outside the colour selection windows or at the limits of the Gaia and WISE sensitivity. In future Euclid data releases, the construction of a more complete quasar sample will be enabled by combining traditional pre-selected candidates with new candidates identified from Euclid photometry and by the SPE pipeline. This composite approach, leveraging the full capability of the Euclid data, will allow for the recovery of quasars with a wider range of colours, luminosities, and host properties.
@@ 6 Conclusions
In this work, we have presented a large sample of 3468 bright quasars covering the redshift range of $0<z\lesssim 4.8$ identified with red-grism spectroscopy of the first Euclid Quick Data Release, including 2686 sources with new spectroscopic identifications relative to existing public compilations. To ensure identification efficiency, we focused on a high-purity input quasar candidate sample of 9214 sources based on GDR3-QSOs (Quaia, CatNorth, and CatSouth) and the AllWISE R90 AGN table. Through a template matching process based on the Pearson correlation function and a visual inspection campaign, we labelled quasars and determine their spectroscopic redshifts. The success rate of the spectroscopic identification is 38 %.
From the identified sample, we estimated an empirical spectroscopic depth of $J_{\scriptscriptstyle\rm E}\lesssim 21.5$ and $H_{\scriptscriptstyle\rm E}\lesssim 21.3$ at the sensitivity of the Wide Field Survey, beyond which the number of securely identified quasars declines sharply. Our investigation of the novel Euclid colour space demonstrates its power in uncovering redder, dust-obscured quasars that may be missed by traditional optical and MIR selections. The synergy between Euclid and WISE photometry promises even greater completeness and diversity in future quasar surveys.
We constructed the first Euclid composite spectrum of bright quasars, covering rest-frame NUV to NIR wavelengths and free from telluric absorption, providing a valuable benchmark for future spectral studies. The spectroscopic bright quasar catalogue of this work, and the composite quasar spectrum, will be available at CDS. In addition, the spectral properties of the quasars will be derived by multi-component fitting (Euclid Collaboration: Calhau et al. in preparation).
We characterised VIS morphologies using single-Sérsic and model-independent (CAS) parameters, supplemented by a deep-learning PSF fraction, $f_{{\scriptscriptstyle\rm PSF}}$. At low redshift ($z<0.5$), obvious host structures are frequently resolved, and a single Sérsic model describes the light profiles of about half of the sources. At intermediate redshift ($0.5<z<2$), the nuclear component dominates, driving Sérsic indices of roughly 90 % of the sources to the upper bound; in this regime $f_{{\scriptscriptstyle\rm PSF}}$ provides a more reliable compactness measure and is used to quantify the balance between nuclear and host emission across the sample.
Finally, we discuss current limitations, including the pre-selection bias and the challenges of host-nucleus decomposition, and outline pathways for improvement with upcoming Euclid data releases and advanced selection methods. The results presented here demonstrate the capability of Euclid slitless spectroscopy in building large quasar samples and provide a basis for extending this approach to the wide-area AGN censuses enabled by forthcoming Euclid data releases.
Data availability.
The full Tables 2 and 5 are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Acknowledgements.
We thank the two anonymous referees for their constructive reports and helpful suggestions, which improved the clarity and quality of this paper. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMIMI, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org/consortium/community/). This work has made use of the Euclid Quick Release Q1 data from the Euclid mission of the European Space Agency (ESA), 2025, https://doi.org/10.57780/esa-2853f3b. This research makes use of ESA Datalabs (datalabs.esa.int), an initiative by ESA’s Data Science and Archives Division in the Science and Operations Department, Directorate of Science. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, and NEOWISE, which is a project of the Jet Propulsion Laboratory/California Institute of Technology. WISE and NEOWISE are funded by the National Aeronautics and Space Administration. This publication makes use of data from the Hyper Suprime-Cam (HSC). This research used data obtained with the Dark Energy Spectroscopic Instrument (DESI). DESI construction and operations is managed by the Lawrence Berkeley National Laboratory. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science and Innovation of Spain (MICINN), and by the DESI Member Institutions: www.desi.lbl.gov/collaborating-institutions. The DESI collaboration is honored to be permitted to conduct scientific research on I’oligam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U.S. National Science Foundation, the U.S. Department of Energy, or any of the listed funding agencies. This research uses services or data provided by the SPectra Analysis and Retrievable Catalog Lab (SPARCL) and the Astro Data Lab, which are both part of the Community Science and Data Center (CSDC) program at NSF National Optical-Infrared Astronomy Research Laboratory. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA), Inc. under a cooperative agreement with the National Science Foundation. This research made use of hips2fits [^8] https://alasky.cds.unistra.fr/hips-image-services/hips2fits, a service provided by CDS. YF is financed by the Dutch Research Council (NWO) grant number OCENW.XL21.XL21.025. KIC acknowledges funding from the Dutch Research Council (NWO) through the award of the Vici Grant VI.C.212.036. VA, JC and FR acknowledge the support from the INAF Large Grant ”AGN & Euclid: a close entanglement” Ob. Fu. 01.05.23.01.14.
Appendix A Sources with initially discrepant redshifts in the DESI comparison
We illustrate the 16 initially discrepant sources in the DESI consistency check, defined as $\left|z_{\mathrm{vi}}-z_{\mathrm{DESI}}\right|/(1+z_{\mathrm{DESI}})>0.15$ among the 454 quasars in common between our Q1 sample and DESI DR1. For each of the 16 sources, we retrieve the DESI DR1 coadded 1D spectrum and the corresponding redshift ($z_{\mathrm{DESI}}$) from the DESI redshift catalogue using the SPectra Analysis and Retrievable Catalog Lab (SPARCL; Juneau et al. 2025) and the Astro Data Lab (Fitzpatrick et al. 2014; Nikutta et al. 2020), and compare them to the Euclid ${\rm RG}_{\scriptscriptstyle\rm E}$ spectrum used in our visual inspection. The assessment focuses on whether the main emission features are mutually consistent between the DESI and Euclid spectra at the proposed redshift solution.
Figure 13 shows the DESI and Euclid spectra together with the template spectrum evaluated at the final $z_{\mathrm{vi}}$. For 11 sources, both DESI and Euclid spectra support our visual redshift solution, and $z_{\mathrm{DESI}}$ is inconsistent with the spectra. For the remaining five sources whose visual redshifts are revised after reinspection, the panel titles list the updated $z_{\mathrm{vi}}$ together with the initial (incorrect) $z_{\mathrm{vi}}$ in parentheses; the updated values are adopted in the final catalogue.
As shown in Fig. 13, most of the 11 sources for which $z_{\mathrm{DESI}}$ is inconsistent with the spectra exhibit an unusually red rest-frame UV continuum. In addition, several of these objects (Euclid Q1 object_id: 2663629828654232723, 2671669943665629138, 2725249569662970807, 2695896801645034348, and 2718727659662564020) show broad absorption troughs, including absorption associated with Mg ii and other metal transitions, consistent with the unusual low-ionisation broad absorption line (LoBAL; see e.g. Boroson and Meyers 1992; Hall et al. 2002; Urrutia et al. 2009) quasars. Because such strongly reddened and LoBAL quasars are relatively rare in the training and template sets used by automated pipelines, and because their red, observed-frame optical continua often have lower S/N than those of typical blue quasars, automated redshift measurements can be less reliable for these objects.
For the remaining five sources, the main causes of the initial misidentifications of $z_{\mathrm{vi}}$ are (i) line-identification degeneracies when only a single prominent feature is present, and (ii) artefacts that complicate the interpretation of the spectrum. In particular, in two cases (object_id: 2689353840663761572, and 2680117565652270034) an emission feature consistent with H $\alpha$ at the correct redshift was initially misidentified as Mg ii, leading to spuriously high redshift estimates ($z_{\mathrm{vi}}>4$). In another case (object_id: 2664764876670636991), an emission feature consistent with H $\beta$ at the correct redshift was initially interpreted as H $\alpha$, leading to an overestimated redshift. Conversely, for object 2675695970669312051, H $\alpha$ was initially misidentified as H $\beta$, leading to an underestimated redshift. Finally, for object 2667540229671259497, the Euclid spectrum is strongly affected by artefacts (anomalous spike and bump), likely related to contamination from a nearby source, and the initial redshift estimate was driven by the spike seen in the Euclid spectrum.
Figure 13: DESI (blue) and Euclid (green) spectra for the 16 objects that initially satisfy $\left|z_{\mathrm{vi}}-z_{\mathrm{DESI}}\right|/(1+z_{\mathrm{DESI}})>0.15$ among the 454 quasars in common with DESI DR1. The dashed red curve shows the quasar template spectrum evaluated at the final $z_{\mathrm{vi}}$ (scaled for display). Vertical dotted lines mark the expected observed wavelengths of common quasar emission lines at the final $z_{\mathrm{vi}}$. Panels outlined in orange denote the five sources for which we revise $z_{\mathrm{vi}}$ after reinspection; for these, the initial (incorrect) $z_{\mathrm{vi}}$ is reported in parentheses after the updated $z_{\mathrm{vi}}$ in the panel title.
Appendix B Description of the spectroscopically identified bright quasar catalogue